Suppose that your long lost relative recently left you an inheritance of $300,000. You decide to invest the
money into an account:
You receive monthly payments while it earns 3.25% compounded monthly. If you
want the payments to last for 40 years, how much should your payments be each month?


Thanks

Answer by Theo(13342)

use this online calculator to find the payments per month you will receive.

http://www.arachnoid.com/lutusp/finance.html

make the following entries into that calculator.
pv = 300,000
fv = 0
np = 40 * 12 = 480
ir = 3.25 / 12 = .270833
select payment at end of time period.
click on pmt.
the calculator should tell you that the monthly payment will be 1117.62 per month.